Required information [The following information applies to the questions displayed below. Nations Capital Fitness, Inc. operates a chain of fitness centers in the Washington, D.C, area. The firms controller is accumulating data to be used in preparing its-annual profit plan for the coming year. The cost behavior pattern of the firms equipment maintenance costs must be determined. The accounting staff has suggested the use of the high-low method to develop an equation, in the form of Y a + bX, for maintenance costs. Data regarding the maintenance hours and costs for last year are as follows: Bours of Maintenance Maintenance Honth January February March April May June July August September October November Service 530 . 470 260 480 350 430 330 440 490 370 370 340 4,860 Costs $ 5,045 4,250 2,750 4,320 3,020 4,160 3,080 3,590 3,970 3,280 3,200 es Total Average $43,745 3,645 405 Rounded 2. Using your answer to requirement (1), what is the variable component of the maintenance cost? (Round your answer to two decimal places.) Next > 89 of14 İİİ
Answer #1

Answer 1:

High and Low values from the data and differences between them are calculated below:

Month Hours of maintenance service Maintenance cost $5,045 $2,750 $2,295 Hiliaiay Low 530 260 270 March Difference

Variable cost per hour of maintenance hour = 2295 / 270 = $8.50

Fixed cost (using data of low) = Maintenance cost - Variable cost = 2750 - 260 * 8.50 = $540

Hence the equation is:

Y = 540 + 8.5 X

Or

Maintenance cost = $540 + $8.50 * Hours of maintenance service

Answer 2:

The variable cost component of the maintenance cost is = $8.50 per maintenance hour